OpenStudy (anonymous):

which of the following is equivalent to 2csc(2x)? a. csc(x) b. sec(x) c. csc(x) sec(x) d. sin(x) cos(x)

OpenStudy (anonymous):

looky looky identity there are only two things you need to recall here, really, apart from the more obvious ones... things like \[\large \csc\theta = \frac1{\sin\theta}\]\[\large \sec\theta = \frac1{\cos\theta}\] just remember that \[\large \sin(2\theta) = 2\sin(\theta)\cos(\theta)\]

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

I'm feeling d., but is there a way to break it down?

OpenStudy (anonymous):

while I'd normally say trust your feelings, this is not the time for that :D I don't actually know ^.^

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

well, sin 2x=2 sin x cos x

OpenStudy (anonymous):

so would csc 2x= 1/2 sin x cos x?

OpenStudy (anonymous):

well naturally ;)

OpenStudy (anonymous):

so the 2csc would equal...

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

multiply the whole expression by 2?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so...

OpenStudy (anonymous):

that would still be 1/2 sin x cos x

OpenStudy (anonymous):

nopity nope :>

OpenStudy (anonymous):

well

OpenStudy (anonymous):

2/4 sin x cos x=1/2 sin x cos x

OpenStudy (anonymous):

don't be silly :P when you multiply a fraction by a number, you only multiply it with the numerator XD

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

so 2/2 sin x cos x?

OpenStudy (anonymous):

sin x cos x?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

or 1/sin x cos x

OpenStudy (anonymous):

what does your heart tell you?

OpenStudy (anonymous):

1/ sin x cos x

OpenStudy (anonymous):

\[\Large \frac1{\sin(x)\cos(x)}= \frac1{\sin(x)}\cdot \frac1{\cos(x)}\]

OpenStudy (anonymous):

which is 1/ sin(x) cos(x)

OpenStudy (anonymous):

which is csc (x)

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

well... it can't be sin(x) cos(X) since our answer is 1/sin(x) cos(x)

OpenStudy (anonymous):

so how is it csc(x)

OpenStudy (anonymous):

because sin(x)= sin(x)(cos x)...

OpenStudy (anonymous):

cause sin 2x= 2 sin x cos x

OpenStudy (anonymous):

that's rich XD

OpenStudy (anonymous):

sin2x = 2sinxcosx granted but I don't see how that means \[\Large \frac1{\sin(x)\cos(x) }= \csc(x)\]

OpenStudy (anonymous):

what I moment...

OpenStudy (anonymous):

the derp moment is coming...

OpenStudy (anonymous):

csc(x) sec(x)?

OpenStudy (anonymous):

cause that would equal 1/sin x times 1/sec x

OpenStudy (anonymous):

and that is the right answer I think

OpenStudy (anonymous):

Of course it is. You really need to practice :/

OpenStudy (anonymous):

but dude, I wish my teacher would be more clear. He is Asian and he has trouble speaking English :(

OpenStudy (anonymous):

I also took trig online

OpenStudy (anonymous):

I would probably learn more if I had a teacher like you

OpenStudy (anonymous):

by the way, could you help me with one more problem?

OpenStudy (anonymous):

aww.. I'm part Asian :|

OpenStudy (anonymous):

and okay, but you're taking the lead with that problem :P

OpenStudy (anonymous):

the lead?

OpenStudy (anonymous):

Only occasional hints from me... ok? that way you learn :)

OpenStudy (anonymous):

ahh, lol

OpenStudy (anonymous):

could you help me with one more problem?

OpenStudy (anonymous):

go ahead

OpenStudy (anonymous):

you got a medal :)